OpenStudy (anonymous):

Check my Answer! Rewrite the fractions with their LCD. \[\tt\frac{ 4 }{ 3x - 6y }{~and~}\frac{ 1 }{ 5x - 10y }\] Out of four answers, I narrowed it down to these two. Which one is the answer? A. \(\Large\tt \frac{ 20 }{ 15(x - 2y)} \frac{ 3 }{ 15(x - 2y)}\) B. \(\Large\tt \frac{ 20(x - 2y)}{ 15(x - 2y)} \frac{ 3(x - 2y) }{ 15(x - 2y)}\)

OpenStudy (anonymous):

@dan815 Which one do you think the answer is?

OpenStudy (anonymous):

OH WAIT> THE ANSWERS ARE INCORRECT> LET ME FIX IT

OpenStudy (dan815):

yep allllllllll wrong

OpenStudy (dan815):

always factor first !!

OpenStudy (anonymous):

OK. There we go. Just refresh your page or leave the post and come back.

OpenStudy (anonymous):

@dan815 ^

OpenStudy (dan815):

the latex is not worrkiing

OpenStudy (anonymous):

;-; It works for me...

OpenStudy (anonymous):

I will send a picture....

OpenStudy (dan815):

\[\tt\frac{ 4 }{ 3x - 6y }{~and~}\frac{ 1 }{ 5x - 10y }\] Out of four answers, I narrowed it down to these two. Which one is the answer? A. \(\Large\tt \frac{ 20 }{ 15(x - 2y)} \frac{ 3 }{ 15(x - 2y)}\) B. \(\Large\tt \frac{ 20(x - 2y)}{ 15(x - 2y)} \frac{ 3(x - 2y) }{ 15(x - 2y)}\)

OpenStudy (dan815):

now i see it :)

OpenStudy (anonymous):

Oh, okay!

OpenStudy (dan815):

okay A is right

OpenStudy (anonymous):

Really? :D

OpenStudy (dan815):

yes you multiple top and bottom by 5 for the left frac and you multiply top and bottom by 3 for the right fraction

OpenStudy (dan815):

GOOD JOB BLOSSOM

OpenStudy (dan815):

A+ sticker

OpenStudy (anonymous):

YAY! Thank you very much!!